OpenStudy (anonymous):

rewrite ^3 sqrt (27x-81) -5 to make it easy to graph using a translation. Describe the graph

OpenStudy (anonymous):

\[ \sqrt[3]{27x-81}-5 \]

OpenStudy (anonymous):

Fill, factor the \(27\) \[ \sqrt[3]{27(x-3)}-5 \]Remember that \(3^3=27\) \[ 3\sqrt[3]{x-3}-5 \]

OpenStudy (anonymous):

If you think about the graph of \(f(x)=\sqrt[3]{x}\)... this is just \[ 3f(x-3)-5 \]Meaning that it is shifted to the right by 3, shifted down by 5, and stretched by a factor of \(3\) vertically